# The spectral genus of an isolated hypersurface singularity and a conjecture relating to the Milnor number

### v1

@@ 7. Relation to other conjectures
7.2.
A family $\mathcal{F}$ of $n$-dimensional isolated hypersurface singularities, defined by germs of holomorphic functions $f\colon(\mathbb{C}^{n+1},0)\to(\mathbb{C},0)$, will be called degenerating if it is endowed with a filter and the function $\mu\colon\mathcal{F}\to\mathbb{R}$ induced by the Milnor number tends to infinity. That is,
@@
    Quasi-homogeneous singularities, with weights $w_{0},\ldots,w_{n}$ converging to 0, proven by Saito [27, §3.7, Example 1].
@@
    Irreducible plane curve singularities, with Puiseux pairs $(k_{1},n_{1}),\ldots,(k_{g},n_{g})$ and $n_{g}\to+\infty$, proven by Saito [27, §3.9, Example 3].
@@
    For convenient, generic singularities, with large Newton polyhedra as in Section 6 above, proven by Almirón–Schulze [1, Theorem 1.1].
7.5.
By applying (7.2) to well-chosen functions, one can derive necessary conditions on a family $\mathcal{F}$ to satisfy K. Saito’s conjecture. We state some of these special features.
@@
7.6.
@@
The second item of the proposition should be compared to Proposition 4.2 and Corollary 6.4, to the effect that (7.4) not only holds in the corresponding cases, but the limit is approached from below. Also, we see that Proposition 4.2 and Corollary 6.4 are compatible with the known cases of K. Saito’s conjecture reviewed in §7.4 above.
@@ 7.7. Hertling’s conjecture
While K. Saito’s conjecture describes the distribution of the spectral numbers of an isolated singularity, Hertling’s conjecture focuses on the variance. We discuss the relationship to our conjecture. Combined with K. Saito’s conjecture, we conclude that Hertling’s conjecture does not seem to imply our conjecture.
7.8.
@@
For plane curve singularities, the weak form of the conjecture is true if $\alpha_{\mu}<2/3$.
@@
Therefore, we see that the criterion provided by the corollary, which requires $\alpha_{\mu}<2/3$, will apply at most for some exceptional singularities of the family $\mathcal{F}$, but not in general.
@@ Acknowledgements
We thank Omid Amini, Jean-Benoît Bost, José Ignacio Burgos Gil, Lars Halvard Halle, Thomas Mordant, Claude Sabbah, Kyoji Saito and Jan Stevens for discussions relating to this article. We are in particular grateful to Ken-Ichi Yoshikawa for many discussions and the hospitality in Kyoto, and for many insights into the analytic torsion. The possibility of the continuity of the analytic torsion on Hilbert schemes was suggested by him. The second author wishes to thank the Department of Mathematics at Chalmers University of Technology and the University of Gothenburg for the hospitality too.

### v3

@@ 7.1. Durfee-type conjecture
7.4.
A family $\mathcal{F}$ of $n$-dimensional isolated hypersurface singularities, defined by germs of holomorphic functions $f\colon(\mathbb{C}^{n+1},0)\to(\mathbb{C},0)$, will be called degenerating if it is endowed with a filter and the function $\mu\colon\mathcal{F}\to\mathbb{R}$, induced by the Milnor number, tends to infinity. That is,
@@
    Quasi-homogeneous singularities, with weights $w_{0},\ldots,w_{n}$ converging to 0, proven by Saito [32, §3.7, Example 1].
@@
    Irreducible plane curve singularities, with Puiseux pairs $(k_{1},n_{1}),\ldots,(k_{g},n_{g})$ and $n_{g}\to+\infty$, proven by Saito [32, §3.9, Example 3]. More generally, Alberich-Carramiñana, Àlvarez Montaner and Gómez-López informed us that they have characterized the sequences of irreducible plane curve singularities which satisfy K. Saito’s conjecture. We refer to [2, Theorem 4.0.1] for the precise statement.
@@
    For convenient, generic singularities, with large Newton polyhedra as in Section 6 above, proven by Almirón–Schulze [3, Theorem 1.1].
7.7.
By applying (7.3) to well-chosen functions, one can derive necessary conditions for a family $\mathcal{F}$ to satisfy K. Saito’s conjecture. We state some of these special features.
@@
We refer the reader to [2, Section 5] for a complementary discussion on the minimal spectral number and K. Saito’s conjecture. In particular, a proof of Proposition 7.3 (2) is also provided therein.
7.8.
@@
The second items of the proposition should be compared to the weak form of the conjecture: under both K. Saito’s and our conjecture, the limit is approached from below. Proposition 4.2 and Corollary 6.4 provide examples of this phenomenon. Also, notice that these statements are compatible with the known cases of K. Saito’s conjecture reviewed in §7.6 above.
@@ 7.9. Hertling’s conjecture
While K. Saito’s conjecture describes the distribution of the spectral numbers of an isolated singularity, Hertling’s conjecture focuses on the variance. We discuss the relationship to our conjecture. Combined with K. Saito’s conjecture, we conclude that Hertling’s conjecture does not seem to trivially imply our strong conjecture for curves.
7.10.
@@
For plane curve singularities, if $\alpha_{\mu}\leq\frac{2}{3}\sqrt{1-\mu^{-1}}$, then the strong form of the conjecture holds.
@@
Therefore, we see that the criterion provided by the corollary, which implies $\limsup_{\mathcal{F}}\alpha_{\mu}\leq 2/3$, will apply at most to some exceptional singularities of the family $\mathcal{F}$, but not in general.
@@ Acknowledgements
We thank Omid Amini, Jean-Benoît Bost, Lars Halvard Halle, Thomas Mordant, Claude Sabbah, Kyoji Saito and Jan Stevens for discussions relating to this article, and Josep Àlvarez Montaner for bringing the article [2] to our attention. We heartily thank José Ignacio Burgos Gil for providing the suspension trick which became Lemma 7.1. We are deeply indebted to Ken-Ichi Yoshikawa for many discussions and the hospitality in Kyoto, and for many insights into the analytic torsion. The possibility of the continuity of the analytic torsion on Hilbert schemes was suggested by him. The second author wishes to thank the Department of Mathematics at Chalmers University of Technology and the University of Gothenburg for the hospitality too.